# Optimal market-neutral currency trading on the cryptocurrency platform

### v2

@@ Abstract
This research proposes a novel arbitrage approach with respect to multivariate pair trading called Optimal Trading Technique (OTT). We introduce the method to selectively form a ‘bucket’ of fiat currencies anchored to cryptocurrency for simultaneously monitoring and exploiting trading opportunities. To handle the quantitative conflicts that arise when receiving multiple trading signals, a bi-objective convex optimization process is designed to cater to the investor’s preference between profitability and risk tolerance. This process includes tunable parameters such as volatility punishment, action thresholds. During our experiments in the cryptocurrency market from 2020 to 2022 when the market was experiencing a vigorous bull-run immediately followed by a bear-run, the OTT realized an annualized profit of 15.49%. We further carried out the experiments in bull, bear, and full-cycle market conditions separately, and found that OTT is capable of achieving stable profit under various market conditions. Apart from the profitability side of the OTT, the arbitrage operation provides a new perspective of trading, which requires no external shorting and never hold intermediate cryptocurrency during the arbitrage period.
@@ 1 Introduction
The secondary market is a modern, efficient form of double auction. Participants with the intention to trade assets bring their acceptable price to bid and ask. The highest bid and lowest ask are to be matched if the price meets. At the macro level, we believe in the efficiency of the secondary market (Fama 1970). Nevertheless, countless economic bubbles bursting in history tells us that investors do not always act rationally. Dale et al. (2005) in the story of South Sea Company depicted a picture of how an orderbook filled with irrational bids and asks could push the price in the wrong direction. Finding out the inefficiencies in an efficient market maintains the equilibrium of market trading. Market directional trading is astute traders estimate a ‘more correct’ pricing for traded instruments and give orders with expectation for the market to revert to being rational. Long stands for traders spend their money to buy the asset for betting on the market to go up and sell later; Short on the opposite means traders borrow the asset and sell to the market to hold money until the market stops dropping to buy back at a cheaper price. Traders that bet in the right direction are rewarded by the market mechanism for correcting the errors. It is a part of the price discovery process. Instead of scrutinizing the market with the help of experienced accountants or financial analysts, we utilize statistics and optimization to dive into applying quantitative trading.
@@
The recent trending cryptocurrency market benefits from the latest technology advancement and global participation. Its volatility, liquidity and 24/7 trading, make the cryptocurrency market an ideal marketplace to conduct algorithmic trading. Since a user can freely trade between cryptocurrencies and fiat currencies within this market, the cryptocurrency market’s volatility comprises the volatility of forex and cryptocurrencies. If we take advantage of the inefficiencies in the emerging cryptocurrency market and capture anomalous movements from the movements of cryptocurrencies, we shall be able to generate profit from the fast-paced market.
@@
The method of trading we focus on in this research is Pair trading. It is one of the most widespread strategies based on the ideology — ‘Law of one price (LOP)’, which suggests that homogeneous items should ideally have the same price everywhere. However, it is often violated because of real world restrictions such as trade frictions and legal regulations (Isard 1977; Miljkovic 1999). That means, though similar assets in a pair should have similar price trends, their pricing independence can sometimes lead to divergence. Betting on the direction of convergence often produces profit. Trading approaches based on LOP are called Arbitrage and are perceived to have more stable profitability. The stability of profiting is favored by investors. Therefore there have been decades of history for investment banks and hedge funds to utilize pair trading as part of their strategies (Gatev et al. 2006; Elliott et al. 2005; Burgess 2000). It is common to find the trading strategy, as the name describes, in pairs. While there has been a history of applying pair trading techniques in real life, we do not yet have many applications for expanding pair trading to multivariate trading. This paper shall take a deep dive into the pair trading technique and introduce an optimization-based technique to ‘pair trade’ multivariate assets and experiment in cryptocurrency market.
@@
This research has multiple novel contributions: ① Unlike classical techniques (Perlin 2007; Galenko et al. 2012; Dunis and Ho 2005), the proposed approach uses optimization to arbitrage multiple pairs simultaneously to the individual asset granularity without grouping assets. ② The presented bi-objective optimization is adjustable to cater to investors’ preferences by the inclusion of tuneable parameters such as risk tolerance, operating thresholds for maximizing return and minimizing the risk. ③ Our arbitrage approach never holds the volatile intermediate cryptocurrency throughout the arbitrage period. By avoiding the hold of cryptocurrency, the position of the trading system consists of only fiat currencies, hence lower risk. ④ The proposed trading approach is designed in a way that no borrowing is required during shorting. All the trading actions are achievable within the basket of the initial principal. Therefore, we save on the cost and hassle of external borrowing. ⑤ Lastly, the proposed technique displays a strong robustness of market-neutrality and profitability in both bull and bear markets.
@@
The remainder of the article is structured as follows. Sections 2 review the most cited works in pair trading and give the crux of the proposed trading idea. Sections 3 introduce the pre-requisite knowledge and the details of the proposed trading methodology, respectively. Simulation algorithm is presented in Section 4. Finally, results and conclusions are presented in Sections 5 and 6, respectively.
@@ 2.1 Pair Trading
Long-Term Capital Management L.P. (LTCM)was a highly-leveraged hedge fund founded by John Meriwether in 1994 operated on bond pair trading. It is famous for its billions of dollars in bailout after the 1997 financial crisis and is known as an early real-life example of pair trading (Slivinski 2009).
@@
The detailed operation of pair trading is rather simple. Figure 1 is a pedagogical demonstration based on the most cited baseline pair trading framework (Gatev et al. 2006), which is also known as Distance Method. The Distance Method of pair trading can be summarized into three steps.
@@
    Exhaustively search for two assets that have the closest price movement in history as y1 and y2. Group y1 and y2 as a trading Pair.
@@
    Difference the price of y1 and y2 as Distance. Set position opening threshold as two times of Distance’s historical standard deviation (presented as the band area with y1). Once the price of y2 surpasses the threshold, open a position by shorting y2 and longing y1. Abide time until the price of y1 and y2 crosses, then close the position by reversing the buy and sell actions from the position opening.
@@
    If y2 goes below the threshold, do the opposite by longing y2 and shorting y1 until the next crossing point. Repeat the practice till end of the trading period.
A critical property of pair trading is that we can achieve Market Neutrality if we carefully maneuver our long and short positions. Market neutrality is the property that the trades are profitable irrespective of the direction the complete market, or the underlying assets being traded move towards. Market neutrality is obtained by a simple principle: if the distance between two correlated assets diverges in the trading period, bet the equivalent quantity on the long side and the short side. The criteria of the divergence can also be versatile. In the work of Chen et al. (2018), the authors form pairs from asset returns instead of their prices, based on the idea that a temporarily weaker return implies a stronger rebounce (and vice versa). From another perspective, both Gatev et al. (2006) and Chen et al. (2018) construct their work based on correlation.
Besides correlation, there is an essential concept in the time-series analysis called cointegration. Cointegration is defined by the co-movement of two time-series data sets and is believed to represent a long-run relationship (Engle and Granger 1991). In the book by Vidyamurthy (2004), he examined assets for cointegration with the Engle-Granger test and bound assets based on not only correlation but also cointegration. His work provided good guidance for cointegration-based pair trading.
A step ahead, there are many studies discussing multivariate pair trading. They adopt the idea of hedging market risk by managing positions and creating ‘artificial pairs’ from scenarios with more than two assets. The technique developed by Perlin (2007) combined assets into an artificial index to form a pair. The main idea is that, instead of finding a pair, he created a synthetic index $P_{it}^{*}$ to anchor a real index $P_{it}$. The synthetic index $P_{it}^{*}$ could be constructed by giving weights ($w$) to $m$ number of assets based on their prices:
@@
We can tune the weights among the participating assets so as to match them close to our asset $P_{it}$ as much as possible. The next step is to arbitrage $P_{it}$ and $P_{it}^{*}$ for pair trading. The methodology of artificial pair exists in cointegration methods as well. Galenko et al. (2012) used a moving window method to pair trade among a group of assets. In the moving window, the authors calculated the cointegration vector $\boldsymbol{b}$ to track $Z_{t}$ from the asset return combined with cointegration value $b$.
@@
Next, assets were partitioned into two categories by the results (positive or negative) of $\boldsymbol{b}$. In the two categories: $L$ (Long) and $S$ (Short), quantified $Z_{t}$ could provide quantifiable recommendations for how much should be invested in each asset. A similar methodology is adopted by Dunis and Ho (2005). They track the EURO STOXX50, and split the outperformers and underperformers into STOXX50+ and STOXX50-. Next, they pair traded STOXX50+ and STOXX50- to achieve ‘double alpha’. Because this method took on non-negligible market risk, therefore it is a weak form of pair trading. We can see that the majority of the existing pair trading techniques are centered around the idea of compounding pairs. In our research, we aim to provide a technique that neither makes an artificial index nor binds a group of assets. With the inclusion of optimization, we can operate directly on all participating assets.
@@ 2.2 Cryptocurrency
The work in Fang et al. (2022) provides a comprehensive survey of the history and recent research about cryptocurrencies. Cryptocurrency was started by the groundbreaking work of Nakamoto (2008) to introduce a decentralized monetary system. The property of decentralization comes from the idea of a peer-to-peer ledger recording system. By incentivizing Bitcoin users to record a full transaction ledger on the user’s own computer, the ledgers are distributed in countless individual nodes instead of a centralized server. The property brings frictionless international participation of Bitcoin. Furthermore, global participation provides excellent liquidity for Bitcoin trades. Not soon after, the success of Bitcoin was followed by imitators. Ethereum (ETH) is the second most famous cryptocurrency in the queue created by Buterin (2014). Ethereum is claimed to be the next generation of blockchain because of its programmable extensibility. In the application, ETH is our quote currency for arbitrage. We choose ETH because of the accessibility of data and high liquidity.
Popular pair trading techniques in Section 2.1 are also introduced in the Cryptocurrency market. The work in Fil and Kristoufek (2020) had a distance method application based on the work of Gatev et al. (2006), which achieved 3% monthly profit in the Cryptocurrency market. The work in Van den Broek and Sharif (2018) formed pairs based on 60-day trading periods to profit on arbitrage opportunities and proved the inefficiency of the cryptocurrency market.
Figure 2: The price of ETH price in multiple currencies
Similar to gold, ETH has a global market in various currencies. An ETH can be traded independently with most major fiat currencies. Trading behavior incurs volatilities, and the volatilities of ETH in different fiat currencies are not always synchronized. The tiny gaps of inefficiencies are the trading opportunities. What makes ETH a better anchor than gold is that ETH as a cryptocurrency is completely fungible and has low international transaction frictions. As we can see in Figure 2. Though the CAD, EUR, GBP and USD follow a similar trend, they do not have a 100% correlation. Our technique expects to capture the minor volatility movement and arbitrage the market inefficiencies.
Though the cryptocurrency system functions in a decentralized manner, cryptocurrency users can deposit their Bitcoin or Ether into a centralized exchange for trading. The exchange works just the same way as stocks or forex exchange. Where you can freely exchange your cryptocurrency for another cryptocurrency or fiat currency, on spot or margin trading. Due to the decentralized nature of participants on the cryptocurrency exchange, it functions 24/7 with liquidity coming from global participation. Due to these properties, there have been experiments on the cryptocurrency exchanges since as early as 2014 (Shah and Zhang 2014; Brandvold et al. 2015; Cheah and Fry 2015).
@@ 3 Methodology
There are two stages in our trading technique. The first stage is a screening of the assets. We select the assets that are correlated and cointegrated with each other to participate in the asset trading bucket. Once the bucket is formed, we monitor the difference between each asset and conduct trades for opening/closing trading signals.
@@ 3.1 Screening
In the screening stage, we need to analyze the historical data so that the assets are suitable for arbitrage. Since the movement of cryptocurrency is impacted by the volatility of both ETH and fiat currencies, if a fiat currency suffers from severe inflation or deflation, the value change of that fiat could mislead our trading signals. Hence, one of the major participating criteria is the volatility of fiats should be small compared to the cryptocurrency. Figure 3 is an illustration of strong fiat depreciation. The difference between AUD and NZD is relatively stable compared to the rapid drop of JPY. If our technique is trading within the period from Mar 2021 to Sept 2022, the depreciating JPY will send a short signal out of its value change with no mean reversion. To select appropriate fiat currencies into the trading bucket, we reviewed the historical trend of fiat currencies to make sure that they have desired statistical relationships.
The first desired statistical property is correlation. We examine the correlation by Pearson’s correlation coefficient (Cohen et al. 2009). Pearson’s correlation coefficient ranges between $[-1,1]$. The higher the absolute value of the coefficient, the stronger association between the two examees. -1 and 1 represent an absolute negative and positive association, and 0 in the between is for no association between datasets.
The second desired statistical property is cointegration. We call the price difference between two assets as spread. In the long term, a widening spread exposes arbitrageurs’ position to bankruptcy risk. That is the incentive to expect the spread between every two assets to have a constant mean and variance. The characteristic of a time-series dataset with a constant mean and variance is called stationary. Cointegration, is the relationship to call the two time-series datasets whose difference is stationary. Cointegration is vital in pair trading because when we form a pair with cointegration in the cryptocurrency market, longing one asset and shorting another is effectively equivalent to holding the price difference. If the spread is stationary with a constant mean, a convergence should always follow a divergence. And the property is called mean-reversion. In this research, the Eagle-Granger test (Engle and Granger 1987) is used to examine the cointegration relationship.
@@ 3.2 Long and Short Trading Signals
After we have selected the assets, we need to set up rules for how and when to trade the multivariate assets. Some questions ahead of us are: How do we ensure market neutrality for multiple assets? How should we define and react to the trading signals? How do we allocate our principal capital?

### v3

# Optimal market-neutral multivariate pair trading on the cryptocurrency platform

@@ Abstract
This research proposes a novel arbitrage approach in multivariate pair trading, termed the Optimal Trading Technique (OTT). We present a method for selectively forming a “bucket” of fiat currencies anchored to cryptocurrency for monitoring and exploiting trading opportunities simultaneously. To address quantitative conflicts from multiple trading signals, a novel bi-objective convex optimization formulation is designed to balance investor preferences between profitability and risk tolerance. This process includes tunable parameters such as volatility penalties and action thresholds. In experiments conducted in the cryptocurrency market from 2020 to 2022, which encompassed a vigorous bull-run followed by a bear-run, the OTT achieved an annualized profit of 15.49%. Further experiments in bull, bear, and full-cycle market conditions demonstrated that OTT consistently achieves stable profits across various market conditions. Additionally, OTT’s arbitrage operation offers a new trading perspective, requiring no external shorting and avoiding intermediate cryptocurrency holdings during the arbitrage period.
@@ 1 Introduction
The secondary market functions as a modern, efficient form of double auction where participants intending to trade assets submit their acceptable bid and ask prices. The highest bid and lowest ask are matched if the prices meet. We believe in the secondary market’s efficiency (Fama 1970), but historical economic bubbles show that investors do not always act rationally. Dale et al. (2005) illustrated how an order book filled with irrational bids and asks could drive prices in the wrong direction. Identifying inefficiencies in an efficient market helps maintain trading equilibrium. Market directional trading involves astute traders estimating a more accurate price for instruments and placing orders expecting the market to revert to rationality. Traders who bet correctly are rewarded by the market mechanism, contributing to price discovery. Quantitative trading, instead of relying on experienced accountants or financial analysts, utilizes statistics and optimization for better price estimations.
@@
The cryptocurrency market is an emerging secondary market. The recent cryptocurrency market trend benefits from technological advancements and global participation. Its volatility, liquidity, and 24/7 trading make it an ideal platform for algorithmic trading. Users can trade between cryptocurrencies and fiat currencies, combining forex and cryptocurrency volatility. Profits can be generated by exploiting inefficiencies and capturing anomalous movements.
@@
Our research focuses on Pair trading, a widespread strategy based on the Law of One Price (Law of one price (LOP)), suggesting that homogeneous items should have the same price everywhere. However, real-world restrictions like trade frictions and legal regulations often violate this principle (Isard 1977; Miljkovic 1999). Similar assets in a pair should have similar price trends, but their pricing independence can lead to divergence, with convergence often yielding profit. Arbitrage-based trading approaches, grounded in LOP, are perceived as having more stable profitability, favored by investors. Investment banks and hedge funds have used pair trading for decades (Gatev et al. 2006; Elliott et al. 2005; Burgess 2000). While pair trading is common, the extension of pair trading to many pairs (multivariate) trading is not well studied. This paper delves into pair trading and introduces an optimization-based technique for multivariate assets, tested in the cryptocurrency market.
@@
This research has several novel contributions: ① Unlike classical techniques (Perlin 2007; Galenko et al. 2012; Dunis and Ho 2005), our approach uses optimization to arbitrage multiple pairs simultaneously to the individual asset level without grouping assets. ② The bi-objective optimization is adjustable to investor preferences, including tunable parameters like risk tolerance and operating thresholds for maximizing return and minimizing risk. ③ Our arbitrage approach avoids holding volatile intermediate cryptocurrency, maintaining positions in fiat currencies to lower risk. ④ No borrowing is required for shorting, as all trading actions are achievable within the initial principal. ⑤ The proposed technique is robust, market-neutral, and profitable in both bull and bear markets.
@@
The remainder of the article is structured as follows. Sections 2 review the most cited works in pair trading and outline the proposed trading idea. Sections 3 introduce the prerequisite knowledge and details of the proposed trading methodology, respectively. Section 4 presents the simulation algorithm. Finally, Sections 5 and 6 present the results and conclusions, respectively.
@@ 2.1 Pair Trading
Long-Term Capital Management L.P. (LTCM)was a highly-leveraged hedge fund founded by John Meriwether in 1994 that operated on bond pair trading. It is famous for its billions of dollars in bailout after the 1997 financial crisis and is known as an early real-life example of pair trading (Slivinski 2009).
@@
The operation of pair trading is straightforward. Figure 1 illustrates the most cited baseline pair trading framework (Gatev et al. 2006), known as the Distance Method. The common trading terminology includes Long positions as buying assets to sell later at a higher price, and Short as borrowing and selling assets, intending to buy them back at a lower price. We can summarize pair trading in five steps:
@@
    Exhaustively search for two assets with the closest historical price movements, y1 and y2. Group y1 and y2 as a Trading Pair.
@@
    Calculate the price difference between y1 and y2 as Distance. Set the position opening threshold as two times the historical standard deviation of the Distance (shown as the band area with y1) and the price crossing as the closing threshold.
@@
    If y2’s price surpasses the threshold, open a position by shorting y2 and longing y1. Wait until y1 and y2 prices cross, then close the position by reversing the buy and sell actions.
  4. 4.
    If y2’s price falls below the threshold, do the opposite: long y2 and short y1 until the next crossing point.
  5. 5.
    Repeat this practice until the end of the trading period.
A critical property of pair trading is achieving Market-Neutrality by betting equivalent amounts on the long and short positions. Since the loss on one position is compensated by the other position, the portfolio value does not move along with the overall market direction. Chen et al. (2018) formed pairs from asset returns instead of prices, suggesting that a temporarily weaker return implies a stronger rebound, and vice versa. Both Gatev et al. (2006) and Chen et al. (2018) based their work on correlation.
Besides correlation, cointegration is a key concept in time-series analysis. Cointegration represents a long-run relationship between two time-series data sets (Engle and Granger 1991). Vidyamurthy (2004) examined assets for cointegration using the Engle-Granger test and combined assets based on correlation and cointegration, providing guidance for cointegration-based pair trading.
Further studies discuss multivariate pair trading, hedging market risk by managing positions and creating artificial pairs from scenarios with more than two assets. Perlin (2007) developed a technique combining assets into an artificial index to form a pair. Instead of finding a pair, he created a synthetic index $P_{it}^{*}$ to anchor a real index $P_{it}$. The synthetic index $P_{it}^{*}$ was constructed by weighting ($w$) $m$ assets based on their prices:
@@
Weights among the participating assets were tuned to match $P_{it}$ as closely as possible. Next, $P_{it}$ and $P_{it}^{*}$ were arbitraged for pair trading. Such an artificially formed pair-trading methodology sometimes also uses cointegration techniques. Galenko et al. (2012) used a moving window method to pair trade among a group of assets, calculating the cointegration vector $\boldsymbol{b}$ to track $Z_{t}$ from the asset return combined with the cointegration value $b$:
@@
Assets were then partitioned into two categories based on $\boldsymbol{b}$ results: $L$ (Long) and $S$ (Short). Quantified $Z_{t}$ provided recommendations on investment amounts for each asset. Dunis and Ho (2005) adopted a similar methodology, tracking the EURO STOXX50, and split outperformers and underperformers into STOXX50+ and STOXX50-. They pair traded STOXX50+ and STOXX50- to achieve “double alpha”, a weak form of pair trading due to non-negligible market risk.
Most existing pair trading techniques focus on compounding pairs. Our research aims to provide a technique that neither creates an artificial index nor binds a group of assets. Using optimization techniques, we can operate directly on all participating assets.
@@ 2.2 Cryptocurrency
The work in Fang et al. (2022) provides a comprehensive survey of the history and recent research on cryptocurrencies. Cryptocurrency began with the groundbreaking work of Nakamoto (2008), introducing a decentralized monetary system. Decentralization stems from a peer-to-peer ledger recording system. By incentivizing Bitcoin users to record a full transaction ledger on their own computers, the ledgers are distributed across countless individual nodes instead of a centralized server. This property enables frictionless international participation in Bitcoin. Furthermore, global participation provides excellent liquidity for Bitcoin trades. Shortly after, Bitcoin’s success led to the creation of imitators. Ethereum (ETH), the second most famous cryptocurrency, was created by Buterin (2014). Ethereum is considered the next generation of blockchain due to its programmable extensibility.
Popular pair trading techniques from Section 2.1 have also been introduced in the cryptocurrency market. Fil and Kristoufek (2020) applied the distance method based on the work of Gatev et al. (2006), achieving a 3% monthly profit in the cryptocurrency market. van_den_broek_cointegration-based_2018 formed pairs based on 60-day trading periods to profit from arbitrage opportunities, demonstrating the inefficiency of the cryptocurrency market.
Figure 2: The price of ETH in multiple currencies
Similar to gold, ETH has a global market against various fiat currencies. ETH can be traded independently with most major fiat currencies. And the unsynchronization of ETH against different fiat currencies is the inefficiencies in the market that provide trading opportunities. There are some advantages of picking ETH as an anchor over gold because it is completely fungible and has low international transaction frictions. As shown in Figure 2, the CAD, EUR, GBP, and USD follow a similar trend but do not have 100% correlation. Our technique aims to capture minor volatility movements and arbitrage market inefficiencies.
Though the cryptocurrency system operates in a decentralized manner, users can deposit their Bitcoin or Ethereum into a centralized exchange for trading. The exchange functions similarly to stock or forex exchanges, allowing users to freely exchange cryptocurrency for another cryptocurrency or fiat currency, either on spot or margin trading. Due to the decentralized nature of participants, cryptocurrency exchanges operate 24/7 with liquidity coming from global participation. These properties have led to experiments on cryptocurrency exchanges since as early as 2014 (Shah and Zhang 2014; Brandvold et al. 2015; Cheah and Fry 2015).
@@ 3 Methodology
There are two stages in our trading technique. The first stage involves screening the assets. We select assets that are correlated and cointegrated with each other to participate in the asset trading bucket. Once the bucket is formed, we monitor the differences between each asset and conduct trades based on opening and closing trading signals.
@@ 3.1 Screening
In the screening stage, we analyze historical data to ensure the assets are suitable for arbitrage. Since movements are impacted by the volatility of both cryptocurrency and fiat currency, severe inflation or deflation in a fiat currency could mislead our trading signals. Therefore, one major criterion is that the volatility of fiat currencies should be small compared to that of cryptocurrency. An example of strong fiat depreciation is illustrated in Figure 3: the difference between AUD and NZD is relatively stable compared to the rapid drop of JPY, which would send a short signal based on its value change without mean reversion. To select appropriate fiat currencies for the trading bucket, we reviewed the historical trends of fiat currencies to ensure they have the desired statistical relationships.
The first desired statistical property is correlation. We examine correlation using Pearson’s correlation coefficient (Cohen et al. 2009), which ranges between $[-1,1]$. The higher the absolute value of the coefficient, the stronger the association between the two variables. A coefficient of -1 represents an absolute negative association, 1 represents an absolute positive association, and 0 indicates no association between datasets.
The second desired statistical property is cointegration. We refer to the price difference between two assets as spread. In the long term, a widening spread exposes arbitrageurs to bankruptcy risk. Therefore, it is essential that the spread between any two assets has a constant mean and variance. A time-series dataset with these characteristics is called stationary. Cointegration describes the relationship between two time-series datasets whose difference is stationary. Cointegration is vital in pair trading because forming a pair with cointegration in the cryptocurrency market means that longing one asset and shorting another effectively holds the price difference. If the spread is stationary with a constant mean, convergence should always follow divergence. This property is called mean-reversion. In this research, the Engle-Granger test (Engle and Granger 1987) is used to examine the cointegration relationship.
@@ 3.2 Long and Short Trading Signals
After selecting the assets, we need to establish rules for trading the multivariate assets. Key questions include: How do we ensure market neutrality for multiple assets? How should we define and react to trading signals? How do we allocate our principal capital?